# Universal Omni-Language Models

### v1

@@ 1 Introduction
  • Although the existing open-source omni-language models have been trained with data in the three modalities, they could barely exceed the performance of random guess accuracy or even hard to follow the instruction when provided with image and audio together.
@@ 2 Related Work
Many studies have been introduced to enable model to perceive complex audio signals and generate text responses based on human instructions. Notable examples in audio perception include SALMONN Tang et al. (2023), BLSPN Wang et al. (2023a), Speech-LLaMAN Wu et al. (2023), and Qwen-Audio Chu et al. (2023b), all of which demonstrate promising capabilities in audio-focused dialogues. Regarding the visual modality, the large visual language model also makes excellent achievements. BLIP2 Li et al. (2023) uses a Q-Former to align the visual knowledge with the textual information during the pre-training phase. LLaVA Liu et al. (2024b) is pre-trained on the GPT-4 generated multimodal language-image instruction-following data and uses a projection to fuse the vision module and the LM. LLaVA-Next Liu et al. (2024a) improves the single-image performance with a cost of an increasing number of image tokens. Based on the framework LLaVA mode, QwenVL Bai et al. (2023), CogVLM Wang et al. (2023b), and Yi-VL Young et al. (2024) all achieve significant success through a huge number of pre-training data. However, most existing multi-modal language models only focus on a certain modal and generate textual responses. While some models can process textual, audio, and visual inputs simultaneously, open-source models in this field exhibit relatively average capabilities compared to their closed-source counterparts. In the following context of this literature, we refer to the omni-language models (OLMs) as those capable of processing at least three different modalities of data simultaneously [^1] At the moment, we target the models concurrently processing with image, audio, and text as a starting point since these are the most well-explored modalities in the field..
@@
The vision-language benchmarks aim to test models’ ability to combine visual and language data in tasks like OCR, spatial awareness, and multimodal reasoning skills. MM-Vet Yu et al. (2023) focuses on visual question answering (VQA), requiring models to interpret visual data and respond to queries. MMBench Liu et al. (2023b) evaluates models via multiple-choice tasks in both Chinese and English, covering diverse domains. MMStar (Chen et al., 2024a) conducts multi-task evaluations to test multimodal fusion capabilities. MMMU Yue et al. (2024) and CMMMU Zhang et al. (2024) assess model performance on complex vision-language tasks, emphasizing sophisticated multimodal reasoning. In addition, there are several audio-understanding benchmarks. Aishell1 Bu et al. (2017), Aishell2 Du et al. (2018), and Librispeech Panayotov et al. (2015) are designed for automatic speech recognition, while ClothoAQA targets audio QA tasks. For automatic audio captioning and vocal sound classification, researchers have curated Clotho Drossos et al. (2020) and VocalSound Gong et al. (2022). However, there is a significant lack of comprehensive understanding benchmarks to assess models’ ability to simultaneously process various modalities, including textual, audio, and visual inputs.
@@ 4 Experiment Settings
We select three groups of MLLM baselines according to the modalities available: i. omni-language models: AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b); ii. vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series team (2024), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024a), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); iii. audio-language models: LTU series Gong et al. (2023), Mu-LLaMA (Liu et al., 2023a), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not categorize them into the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024..
@@
Table 2: Overall Omini-Undesratnding Results on Baseline Omni-Language Models. The overall, image-ablated and audio-ablated results on all samples are put at the first three columns. The accuracy calculated by different audio types the last three columns.
Accuracy $\uparrow$ All Audio Types Speech Sound Event Music
Input Context Image & Audio Audio Image Image & Audio
AnyGPT (7B) 2.71% 2.36% 1.23% 3.11% 1.89% 1.89%
video-SALMONN (13B) 11.30% 11.56% 11.38% 13.49% 9.06% 0.94%
UnifiedIO2-large (1.1B) 22.68% 24.69% 24.52% 21.92% 23.02% 27.36%
UnifiedIO2-xlarge (3.2B) 20.40% 24.78% 24.34% 22.96% 15.09% 15.09%
UnifiedIO2-xxlarge (6.8B) 23.29% 27.06% 25.04% 25.29% 19.62% 17.92%
Gemini-1.5-Pro 47.56% 38.53% 34.68% 47.90% 46.36% 48.11%
Reka-core-20240501 36.10% 35.07% 34.39% 38.06% 29.81% 37.74%
5 Findings
5.1 Results on Existing Omni-Language Models
Table 2 demonstrates that open-source omni-language model (OLM) baselines barely surpass random guessing accuracy across various settings. The lack of significant performance improvement in models with larger text backbones suggests that current open OLMs struggle to effectively interpret multimodal contexts and comprehend relationships between elements.
@@ 5.1 Results on Existing Omni-Language Models
InternVL-2-2B 30.04% 19.61% 23.29%
InternVL-2-8B 35.38% 24.69% 27.76%
InternVL-2-26B 34.68% 22.77% 26.80%
InternVL-2-40B 35.03% 24.34% 26.53%
Qwen2-VL-Chat-2B 30.21% 22.33% 30.74%
Qwen2-VL-Chat-7B 33.45% 24.96% 30.30%
Deepseek-VL-Chat-7B 27.67% 20.75% 23.12%
Idefics2-8B 7.09% 11.65% 1.66%
Mantis-Idefics-8B 4.29% 2.36% 0.26%
LLaVA-OneVision-0.5B 33.01% 30.56% 30.65%
LLaVA-OneVision-7B 33.27% 25.48% 23.29%
Cambrian-8B 30.65% 23.56% 25.57%
Cambrian-13B 35.64% 25.83% 31.26%
Cambrian-34B 28.46% 20.14% 22.24%
XComposer2-4KHD (7B) 30.91% 20.75% 24.87%
@@
UnifiedIO2-large (1.1B) 30.56% 31.00% 24.52%
UnifiedIO2-xlarge (3.2B) 27.32% 32.57% 24.34%
UnifiedIO2-xxlarge (6.8B) 33.63% 27.85% 25.04%
@@
UnifiedIO2-large (1.1B) 27.85% 24.69% 28.81% 27.50% 30.19% 24.53%
UnifiedIO2-xlarge (3.2B) 26.80% 24.78% 29.07% 28.53% 24.53% 19.81%
UnifiedIO2-xxlarge (6.8B) 28.46% 27.06% 26.44% 27.89% 33.96% 18.87%
@@ 5.2 Textual Approximation on Images and Audios
We select the UnifiedIO-2 series OLMs to conduct the textual approximation experiments due to their relatively robust performances in the vanilla OmniBench evaluation setting suggested in Table 2. Compared with the vanilla setting, UnifiedIO-2 models shows different levels of performance gains in this settings, i.e. averagely $8.38\%$, $5.58\%$ and $9.4\%$ when replacing the audio, image, and both of them with corresponding texts. This indicates the shortcoming of existing OLMs on modeling the image and audio (especially the latter), and a defect of interpreting these multimodal information correctly in the instruction context.
@@
InternVL-2-8B 32.84% 24.69% 27.67%
InternVL-2-40B 30.30% 24.34% 25.48%
Cambrian-13B 32.57% 25.83% 27.15%
@@
UnifiedIO2-large (1.1B) 29.77% 31.00% 28.81%
UnifiedIO2-xlarge (3.2B) 31.70% 32.57% 29.07%
UnifiedIO2-xxlarge (6.8B) 33.10% 27.85% 26.44%

### v2

@@ 1 Introduction
The social impact of these MLLMs is far-reaching, providing transformative capabilities for a variety of domains. In healthcare, VLM and ALM can contribute to diagnosing Wu et al. (2023a); Hemdan et al. (2023). The integration of all modalities is expected to significantly improve diagnostic accuracy and patient interaction, making healthcare more accessible and efficient. In urban environments, ALM can contribute to improving safety and traffic management by incorporating urban sound event detection during autonomous driving, such as recognising audio of emergency vehicles to supplementary visual and radar modality Sun et al. (2021). In addition, audio contributes to biodiversity monitoring Terenzi et al. (2021); Liang et al. (2024a) and can be greatly enhanced by MLLM’s ability to analyse both audio and video from a variety of sensors. Finally, it may help robotics or LLM agents to provide better human-computer/robotic interaction (HCI/HRI) service in day-to-day life Liang et al. (2024b); Su et al. (2023).
@@
  • Although the existing open-source omni-language models have been trained with data in the three modalities, most of them can surpassthe performance of random guess accuracy but sometimes hard to follow the instruction when provided with image and audio together in certain cases.
@@ 2 Related Work
Whisper is an ASR model Radford et al. (2022) and is used for a speech encoder in SALMONN and LTU. BEATs (Chen et al., 2022) is a general audio encoder used in SALMONN, MERT is an acoustic music encoder (Li et al., 2023b) used in Mu-llama and MusiLingo, CLAP is a music-text encoder (Wu et al., 2023c) used in audio-flamingo,
Many studies have been introduced to enable model to perceive complex audio signals and generate text responses based on human instructions. Notable examples in audio perception include SALMONN (Tang et al., 2023), BLSPN (Wang et al., 2023a), Speech-LLaMAN (Wu et al., 2023b), and Qwen-Audio (Chu et al., 2023b), all of which demonstrate promising capabilities in audio-focused dialogues. Regarding the visual modality, the large visual language models, often equipped with pre-trained image encoders (Dosovitskiy, 2020; Touvron et al., 2020; Liu et al., 2021; Radford et al., 2021; Zhai et al., 2023), also makes excellent achievements. BLIP2 (Li et al., 2023a) uses a Q-Former to align the visual knowledge with the textual information during the pre-training phase. LLaVA (Liu et al., 2024b) is pre-trained on the GPT-4 generated multimodal language-image instruction-following data and uses a projection to fuse the vision module and the LM. LLaVA-Next (Liu et al., 2024a) improves the single-image performance with a cost of an increasing number of image tokens. Based on the framework LLaVA mode, QwenVL (Bai et al., 2023), CogVLM (Wang et al., 2023b), and Yi-VL (Young et al., 2024) all achieve significant success through a huge number of pre-training data. However, most existing multi-modal language models only focus on a certain modal and generate textual responses. While some models can process textual, audio, and visual inputs simultaneously, open-source models in this field exhibit relatively average capabilities compared to their closed-source counterparts. In the following context of this literature, we refer to the omni-language models (OLMs) as those capable of processing at least three different modalities of data simultaneously [^1] At the moment, we target the models concurrently processing with image, audio, and text as a starting point since these are the most well-explored modalities in the field..
@@
The vision-language benchmarks aim to test models’ ability to combine visual and language data in tasks like OCR, spatial awareness, multimodal information retrieval (e.g. SciMMIR (Wu et al., 2024a)), and multimodal reasoning skills. MM-Vet (Yu et al., 2023) focuses on visual question answering (VQA), requiring models to interpret visual data and respond to queries. MMBench (Liu et al., 2023b) evaluates models via multiple-choice tasks in both Chinese and English, covering diverse domains. MMStar (Chen et al., 2024a) conducts multi-task evaluations to test multimodal fusion capabilities. MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024) assess model performance on complex vision-language tasks, emphasizing sophisticated multimodal reasoning. MMRA Wu et al. (2024b) is designed to evaluate the models’ multi-image relational association capability. In addition, there are several audio-understanding benchmarks. Aishell1 (Bu et al., 2017), Aishell2 (Du et al., 2018), and Librispeech (Panayotov et al., 2015) are designed for automatic speech recognition, while ClothoAQA targets audio QA tasks. For automatic audio captioning and vocal sound classification, researchers have curated Clotho (Drossos et al., 2020) and VocalSound (Gong et al., 2022). However, there is a significant lack of comprehensive understanding benchmarks to assess models’ ability to simultaneously process various modalities, including textual, audio, and visual inputs.
@@ 4 Experiment Settings
We select three groups of MLLM baselines according to the modalities available: i. omni-language models: AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b); ii. vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series (team, 2024), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024a), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); iii. audio-language models: LTU series (Gong et al., 2023), Mu-LLaMA (Liu et al., 2023a), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not categorize them into the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024..
@@
Table 2: Overall Omini-Undesratnding Results on Baseline Omni-Language Models. The overall, image-ablated and audio-ablated results on all samples are put at the first three columns.
Input Context Image & Audio Audio Image
AnyGPT (7B) 18.04% 16.20% 20.05%
video-SALMONN (13B) 35.64% 35.90% 34.94%
UnifiedIO2-large (1.1B) 27.06% 29.07% 29.07%
UnifiedIO2-xlarge (3.2B) 38.00% 31.17% 34.76%
UnifiedIO2-xxlarge (6.8B) 33.98% 32.49% 33.45%
Gemini-1.5-Pro 47.56% 38.53% 34.68%
Reka-core-20240501 36.10% 35.07% 34.39%
Table 3: OLM Baselines Overall Results Grouped by Audio Type and Task Category. The accuracy calculated by different audio types are at the first three columns and the accuracy accross tasks categories are placed at the rest columns.
Model Speech Sound Event Music
AnyGPT (7B) 17.77% 20.75% 13.21%
video-SALMONN (13B) 34.11% 31.70% 56.60%
UnifiedIO2-large (1.1B) 25.94% 29.06% 30.19%
UnifiedIO2-xlarge (3.2B) 39.56% 36.98% 29.25%
UnifiedIO2-xxlarge (6.8B) 34.24% 36.98% 24.53%
Gemini-1.5-Pro 47.90% 46.36% 48.11%
Reka-core-20240501 38.06% 29.81% 37.74%
Model Action & Activity Story Description Plot Inference Object Identification & Description Contextual & Environmental Identity & Relationship Text & Symbols Count & Quantity
AnyGPT (7B) 19.52% 16.52% 14.77% 22.27% 15.60% 21.88% 12.00% 33.33%
video-SALMONN (13B) 31.47% 28.26% 25.74% 62.56% 36.88% 37.50% 20.00% 6.67%
UnifiedIO2-large (1.1B) 29.88% 20.87% 31.65% 30.81% 23.40% 18.75% 24.00% 6.67%
UnifiedIO2-xlarge (3.2B) 32.27% 33.48% 31.65% 63.03% 34.04% 34.38% 24.00% 20.00%
UnifiedIO2-xxlarge (6.8B) 32.27% 29.13% 29.96% 48.82% 34.75% 25.00% 8.00% 46.67%
@@
InternVL-2-2B 42.29% 27.32% 28.11%
InternVL-2-8B 50.79% 33.63% 33.36%
InternVL-2-26B 51.75% 31.87% 33.89%
InternVL-2-40B 54.29% 31.96% 34.76%
Qwen2-VL-Chat-2B 42.47% 31.44% 38.09%
Qwen2-VL-Chat-7B 48.60% 32.05% 36.87%
Deepseek-VL-Chat-7B 39.67% 29.51% 26.27%
Idefics2-8B 45.10% 32.31% 34.41%
Mantis-Idefics-8B 46.15% 36.43% 32.57%
LLaVA-OneVision-0.5B 38.00% 31.79% 31.17%
LLaVA-OneVision-7B 47.02% 31.70% 29.68%
Cambrian-8B 42.12% 31.35% 32.22%
Cambrian-13B 45.01% 31.96% 33.98%
Cambrian-34B 46.76% 30.12% 33.01%
XComposer2-4KHD (7B) 43.96% 29.25% 30.65%
@@
UnifiedIO2-large (1.1B) 34.33% 31.96% 29.07%
UnifiedIO2-xlarge (3.2B) 43.17% 34.50% 34.76%
UnifiedIO2-xxlarge (6.8B) 40.81% 29.77% 33.45%
@@
UnifiedIO2-large (1.1B) 29.16% 29.07% 29.33% 28.40% 32.45% 26.42%
UnifiedIO2-xlarge (3.2B) 32.22% 31.17% 30.21% 32.43% 32.45% 30.19%
UnifiedIO2-xxlarge (6.8B) 32.05% 32.49% 27.15% 31.13% 38.87% 21.70%
5 Findings
5.1 Results on Omni-Language Models
Table 2 demonstrates that open-source omni-language model (OLM) baselines surpass random guessing accuracy across various settings. Notably, the UnifiedIO2 series demonstrates inconsistent performance scaling with model size, indicating challenges in effectively leveraging increased capacity for multimodal understanding.
Breakdown Results.
We present the breakdown of the performance of open-source omni-language model baselines across different audio types and task categories in the OmniBench evaluation. The results reveal inconsistent performance patterns across audio types, with some models showing higher accuracy on sound events or music compared to speech. Across task categories, models tend to perform better on object identification and description tasks, while struggling with more reasoning tasks such as plot inference and story description. These findings highlight the current limitations of OLMs in integrating information across modalities.
@@ 5.2 Textual Approximation on Images and Audios
We select the UnifiedIO-2 series OLMs to conduct the textual approximation experiments due to their relatively robust performances in the vanilla OmniBench evaluation setting suggested in Table 2. Compared with the vanilla setting, all three UnifiedIO-2 models show performance gains, averagely at $6.42\%$, in the audio replacement setting and average performance drops in the replaced-image ($1.87\%$) and both-repaced settings ($0.12\%$). This indicates the shortcoming of existing OLMs on modeling the audio on the one hand, and there is potentially more noise in the generated image captions compared to the human-written audio transcripts on the other hand.
@@
InternVL-2-8B 47.55% 33.63% 29.86%
InternVL-2-40B 47.20% 31.96% 30.65%
Cambrian-13B 43.08% 231.96% 29.16%
@@
UnifiedIO2-large (1.1B) 30.74% 31.96% 29.33%
UnifiedIO2-xlarge (3.2B) 33.80% 34.50% 30.21%
UnifiedIO2-xxlarge (6.8B) 34.15% 29.77% 27.15%